OpenStudy (anonymous):

what are the zeros in (8x+3)/(11-x)????

OpenStudy (anonymous):

equate the equation to zero and see what you get

OpenStudy (anonymous):

i know i dont know how

OpenStudy (anonymous):

you dont know how to equate the eqution to 0?

OpenStudy (anonymous):

idk where to start! like whats the first step

OpenStudy (anonymous):

8x+3/11-x = 0 that is the first step. solve for x

OpenStudy (anonymous):

i know what do i pull over to the zero side first to get x by itself

OpenStudy (anonymous):

multiply both sides by 11-x and post what you get

OpenStudy (anonymous):

8x+3=11-x....?

OpenStudy (anonymous):

what is 0 times 10?

OpenStudy (anonymous):

o

OpenStudy (anonymous):

0**

OpenStudy (anonymous):

then why is 11-x times 0 any different?

OpenStudy (anonymous):

oh yeah duh. okay so then its just 8x+3=0

OpenStudy (anonymous):

now solve for x

OpenStudy (anonymous):

oh thats easy! duh. thanks!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

so 8/3=x

OpenStudy (anonymous):

no. you made a mistake. try again

OpenStudy (anonymous):

8/-3 is what i meant to say

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

ugh! -3/8!!!

OpenStudy (anonymous):

yes. correct

OpenStudy (anonymous):

thank you! ugh i swear im not this dumb. haha idk whats wrong with me. thank you

OpenStudy (anonymous):

lol you are welco,me